OpenStudy (anonymous):

Karen has 18 dimes and quarters in her pocket that total $3.00. Which of the following pairs of equations would you use to find out how many of each coin she has? Let d represent dimes and q represent quarters. A. d + q = 3.00 .10d + .25q = 18 B. d + q = 18 d + q = 3.00 C. d + q = 18 .10d + .25q = 3.00 D. .10d + q = 18 .10d + .25q = 3.00

OpenStudy (anonymous):

I got D.

OpenStudy (austinl):

That is not what I would use.

OpenStudy (calculusxy):

I think that it is B because you need to first let people know that you have to add the amount of dimes and quarters together. Since you already know the amount of dimes, you just need to obtain how many quarters will you need. Then you can add the dimes and quarters together to get $3.00.

OpenStudy (anonymous):

Okay thanks guys!

OpenStudy (calculusxy):

Your welcome :)

OpenStudy (austinl):

@calculusxy That is not what I would use either.

OpenStudy (austinl):

You have so many dimes at a value plus so many quarters at a value that add up to $3. What would that be?

OpenStudy (anonymous):

C would be a better choice then?

OpenStudy (austinl):

Indeed.

OpenStudy (calculusxy):

But what doesn't make sense to me is "18 .10d ".

OpenStudy (anonymous):

I am confuses about that as well.

OpenStudy (austinl):

It is the formatting. 18 is the end of one equation. .10d is the beginning of the other.

OpenStudy (calculusxy):

But you don't have 10 dimes.

OpenStudy (austinl):

\(d + q = 18\) \(.10d + .25q = 3.00\)

OpenStudy (austinl):

A dime is \(.10\) or \(\dfrac{1}{10}\) of a dollar.

OpenStudy (austinl):

This equation required a knowledge of the American currency system.

OpenStudy (calculusxy):

@austinL It says specifically 18 dimes not d+q=18. I hope you understand.

OpenStudy (austinl):

18 dimes \(\color{red}{\text{AND}}\) quarters...... poorly written problem.

OpenStudy (austinl):

For the record, it would be 8 Quarters and 10 Dimes.

OpenStudy (calculusxy):

Yes sorry for my mistake